A pulsed laser has an average output power of 12W. Each pulse consists of light at
wavelength 400nm and has 16ms duration. Compute the number of photons that are emitted
in a single pulse?
1
Expert's answer
2020-06-08T10:25:17-0400

The total energy of pulse


"E=P\tau=N\frac{hc}{\lambda}"

Hence


"N=\frac{P\tau\lambda}{hc}\\
=\frac{12\times 16\times 10^{-3}\times 400\times 10^{-9} }{6.63\times 10^{-34}\times 3\times 10^8}=3.86\times 10^{17}"